OpenStudy (anonymous):

You have been told that a certain house is 164 m2 in area. How much is this in square feet?

OpenStudy (aaronq):

Find the side length \(Area=side^2\rightarrow side=\sqrt A=\sqrt{164~m^2}=12.8~m\)  1 ft =0.3048 m Convert side of 12.8 m to feet \(12.8~m*\dfrac{1~ft}{0.3048~m}=41.9947~ft\approx 42~ft\) \(Area=side^2=(42~ft)^2=1764~ft\)

OpenStudy (aaronq):

Correction: \(1764~ft^2\)**